OpenStudy (anonymous):

how to find out the area of a 4 1/3 base and 3 2/5 height parallelogram

OpenStudy (kropot72):

If the height is the vertical height the area is found by multiplying the base by the height: \[\large Area=\frac{13}{3}\times\frac{17}{5}=(you\ can\ calculate)\ sq.\ units\]

OpenStudy (aum):

Area A of a parallelogram = base x height. \(A = \Large 4\frac 13 \times 3 \frac 25 = \frac{13}{3} \times \frac{17}{5} = \frac{13\times 17}{3 \times 5} = \frac{?}{?}\)

OpenStudy (anonymous):

uhm so 13 x 17 and 3 x 5 equals area 14.73????

OpenStudy (kropot72):

The area is 14.73 square units, to two decimal places.

OpenStudy (anonymous):

thank you :D i cant believe how simple it was but i didnt get it

OpenStudy (kropot72):

You're welcome :) note: If the slant height is given instead of the vertical height, an internal angle must be given. Then it necessary to use simple trigonometry to find the vertical height.

OpenStudy (aum):

Since the original dimensions were given as a mixed number I would probably leave the answer in mixed number too instead of decimals: \(\Large \frac{13\times 17}{3 \times 5} = \frac{221}{15} = 14\frac{11}{15}\)